### v2

@@ 4.2 Benefiting from communication
Thus, the sender benefits from communication if and only if there exists a top equilibrium that the sender strictly prefers to the (sender-preferred) no-communication outcome. The price range in Theorem 2 can be visualized in Figure 1, where $N-1=3$.
@@
Next, suppose $\underaccent{\bar}{\nu}<p\leq\bar{\nu}$. Over this range, without communication, the receiver buys if and only if the checked attribute is good. Recall from Theorem 1 that $\underaccent{\bar}{\nu}<\bar{p}_{N-1}<\bar{\nu}$. If $p\leq\bar{p}_{N-1}$, then the top-$(N-1)$ equilibrium exists. Thus, the sender strictly benefits from communication by guiding the receiver to check attributes that are better than average. If $p>\bar{p}_{N-1}$, then no top equilibrium exists. In principle, there could be other equilibria that the sender prefers to the no-communication outcome. Our proof rules this out, using the bounds in Lemma 1.
Finally, if $p>\bar{\nu}$, then without communication, the receiver never buys. In this case, we use Lemma 1 to show that there is no equilibrium in which the receiver buys with positive probability. The sender’s incentive constraints are critical to this conclusion. For prices $p$ slightly above $\bar{\nu}$, the receiver would be willing to buy upon seeing that one of the worst attributes was good. But in equilibrium, the sender cannot induce the receiver to check a below-average attribute. The sender would strictly prefer the receiver, while thinking he was checking one of the worst attributes, to actually check one of the best.
@@ 4.3 Optimality of top equilibria
By Theorem 1, the top-$k$ equilibrium is the sender-preferred top equilibrium if and only if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$. Therefore, Theorem 3 implies that the sender-preferred top equilibrium is Pareto efficient. To prove Theorem 3, we show that if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$, then no vector of buying probabilities that satisfies the bounds in Lemma 1 can Pareto dominate the top-$k$ equilibrium. If $p=\bar{p}_{k}$, then sender optimality follows from Pareto efficiency because the top-$k$ equilibrium gives the receiver the lowest possible payoff among all equilibria—the receiver’s payoff from the top-$k$ equilibrium is the same as his payoff from checking a random attribute and buying if and only if it is good. The prices $\bar{p}_{1},\ldots,\bar{p}_{N-1}$ can be visualized in Figure 1 as the right endpoints of the equilibrium price intervals.
Up to this point, we have assumed that the price $p$ is exogenous. Next, we assume that the sender can choose the price (before observing the state). Thus, the sender can choose the price to maximize her revenue in the resulting equilibrium. We characterize which prices can be optimal. To state this result, it is convenient to extend our definition of the top-$k$ equilibrium to the edge cases $k=0$ and $k=N$. Under the top-00 0 strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and then always buys. Under the top-$N$ strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and buys if and only if the checked attribute is good. [^8] In each case, there are multiple strategy profiles with this property, but all such strategy profiles are payoff equivalent. Let $\bar{p}_{0}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N}=\bar{\nu}$. For $k=0$ and $k=N$, it is clear that $\bar{p}_{k}$ is the highest price at which the top-$k$ strategy profile is an equilibrium.
@@
We illustrate the intuition for Theorem 4 graphically. Recall that $Q$ denotes the set of buying probability vectors $q=(q_{0},\ldots,q_{N})\in[0,1]^{N+1}$ satisfying the bounds in Lemma 1. The sender’s and receiver’s payoffs from a vector $q\in Q$ are determined by two statistics: the ex-ante buying probability $x(q)$ and the expected consumption utility $y(q)$. In particular, the sender’s payoff is $px(q)$ and the receiver’s payoff is $y(q)-px(q)$. To compute these statistics, we introduce additional notation. For $j=0,\ldots,N$, let $\pi_{j}$ denote the probability $\mathbf{P}(\lvert\theta\rvert=j)$, and let $v_{j}$ denote the consumption utility $v(\theta)$ in any state $\theta$ with $\lvert\theta\rvert=j$ [^9] This is well-defined because $v$ is symmetric across the components. We have
@@
The set $\mathcal{X}$ is more convenient to analyze than the set $Q$ because it lies in $\mathbf{R}^{2}$ rather than $\mathbf{R}^{N+1}$. Figure 2 plots the set $\mathcal{X}$ in the numerical example with $N=4$ used in Figure 1. The upper envelope of $\mathcal{X}$ is highlighted. The extreme points of this upper envelope correspond to the top-$k$ equilibria for $k=0,\ldots,N$ and also the origin, as we show in Lemma 3 (Section A.4). At price $p$, the receiver’s indifference curves in $\mathcal{X}$ are parallel lines with slope $p$. For $k=0,\ldots,N-1$, at price $\bar{p}_{k}$ the receiver is indifferent between the top-$k$ equilibrium and deviating by checking a random attribute and buying if and only if that checked attribute is good. Therefore, $\bar{p}_{k}$ is the slope of the line segment between top-$k$ equilibrium vector and the top-$N$ equilibrium vector. Using the characterization of the upper envelope of $\mathcal{X}$, we compute for each price $p$ an upper bound $V(p)$ on the sender’s revenue from all equilibria at price $p$. For $k=0,\ldots,N$, the bound $V(\bar{p}_{k})$ equals the sender’s revenue from the top-$k$ equilibrium at price $\bar{p}_{k}$. To prove Theorem 4, it suffices to show that the function $V$ is globally maximized at some price in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. The function $V$ has kinks at these prices. We show that $V$ is convex between consecutive kinks, and hence the maximum must be achieved at one of the kinks.
The optimal price depends on the state distribution and the utility function $v$. Consider the simple example with $N=2$ and $v(\theta)=\lvert\theta\rvert$. It can be verified that the top-11 11 equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-22 22 equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. For the calculations, see Section A.8. When $\lvert\theta\rvert=1$, the receiver buys with certainty in the top-11 11 equilibrium but only with probability $1/2$ in the top-22 22 equilibrium. Thus, as $\pi_{1}/\pi_{2}$ increases, the buying probability ratio between the top-11 11 equilibrium and the top-22 22 equilibrium increases. The price ratio $\bar{p}_{1}/\bar{p}_{2}$ also increases because the receiver, upon seeing that a randomly chosen attribute is 11 11, assigns lower probability to the event $\lvert\theta\rvert=2$.
@@ 5 Checking multiple attributes
We now extend the notion of a top-$k$ equilibrium to this setting. The top-$k$ messaging strategy for the sender is defined as in the baseline model: the sender’s message indicates the top $k$ attributes, with ties broken uniformly. The receiver’s top-$k$ strategy is as follows. The receiver checks $n\wedge k$ of the recommended attributes. Then the receiver buys if and only if the number of good attributes among those checked is at least some threshold, which is between 11 11 and $n\wedge k$ [^10] Technically, we have assumed that the receiver must check $n$ attributes. Thus, the receiver can check $n-(n\wedge k)$ of the unrecommended attributes, but their values do not affect his buying probability. We say that the top-$k$ equilibrium exists if such a strategy profile is an equilibrium. [^11] In order to speak of the top-$k$ equilibrium, we assume that if there are multiple such thresholds that make the receiver’s strategy a best response, the smallest threshold is chosen.
@@
When the receiver can check multiple attributes, the set of deviations is much richer. To prove Theorem 5, the trick is to compute the receiver’s best response in a hypothetical specification in which the receiver checks $n\wedge k$ of the $k$ recommended attributes and $n\wedge(N-k)$ of the $N-k$ unrecommended attributes. For simplicity, suppose here that $n\leq k$ and $n\leq N-k$. The random vector $(T_{n}^{k},B_{n}^{N-k})$ can only take the values
@@
We have assumed that the receiver selects all $n$ attributes before seeing any of their values. If the receiver could instead check the attributes adaptively, based on the realizations of previously checked attributes, then our proof of Theorem 5 would still go through. Checking $n$ attributes adaptively is still less informative than checking $n\wedge k$ of the recommended attributes and $n\wedge(N-k)$ of the unrecommended attributes.
@@ A.1 Favorability
is weakly increasing in $j$ over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$; we use the convention that $a/0=\infty$ for $a>0$. If, further, the ratio in (3) is not constant over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$, then we say that $E$ is strictly more favorable than $E^{\prime}$. [^12] Our terminology is inspired by Milgrom (1981). Unlike in his definition, we do not assume that the events have full support. Our formal definition mirrors the definition of the likelihood ratio order in Shaked and Shanthikumar (2007, 1.C.1, p. 42). If $E$ is strictly more favorable than $E^{\prime}$, then $\operatorname{\mathbf{E}}[v(\theta)\mid E]>\operatorname{\mathbf{E}}[v(\theta)\mid E^{\prime}]$, by standard results about the likelihood ratio order.131313 If $E$ is weakly more favorable than $E^{\prime}$, then the distribution of $\lvert\theta\rvert$, conditional on $E$, likelihood ratio dominates the distribution of $\lvert\theta\rvert$, conditional on $E^{\prime}$, in the sense of Shaked and Shanthikumar (2007, 1.C.1, p. 42). Moreover, likelihood ratio dominance implies first-order stochastic dominance (Shaked and Shanthikumar, 2007, Theorem 1.C.1, p. 43).
@@ A.2 Proof of Theorem 1
In Section A.9, we prove a more general version of Theorem 5 that also covers the case $n=1$. Moreover, we can apply Lemma 5 (Section A.9) with $k=N$ and $n=1$ to conclude that $\underaccent{\bar}{p}_{N-1,1}<\underaccent{\bar}{p}_{N,1}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N-1,1}<\bar{p}_{N,1}=\bar{\nu}$. Thus, it remains to prove that $\underaccent{\bar}{\nu}<\bar{p}_{1}$. We prove that the event $(T^{1},B^{N-1})=(1,0)$ is strictly more favorable than the event $\theta_{1}=0$; see Section A.1 for the definition. The event $(T^{1},B^{N-1})=(1,0)$ has support $\{1,\ldots,N-1\}$. The event $\theta_{1}=0$ has support $\{0,\ldots,N-1\}$. We have
@@ A.4 Preliminaries for Theorems 24
For $k=0,\ldots,N+1$ let $(x_{k},y_{k})=(x(q^{k}),y(q^{k}))$. Since the map $q\mapsto(x(q),y(q))$ is linear, the upper envelope of $\mathcal{X}$ can alternatively be expressed as $\cup_{k=0}^{N}[(x_{k},y_{k}),(x_{k+1},y_{k+1})]$. Figure 2 illustrates these $N+1$ line segments in an example with $N=4$.
@@ A.5 Proof of Theorem 2
Recall from Section A.4 the definitions of the vectors $\Delta^{0},\ldots,\Delta^{N}$ and the functions $u_{0},\ldots,u_{N}$. We have
@@
For each $q\in Q$, each coefficient on $u_{k}(p)$ in the summation is nonnnegative. Fix a price $p$ satisfying $p>\bar{p}_{N-1}$. By Lemma 2.iiiii, we have $u_{k}(p)<0$ for $k=1,\ldots,N-1$. By direct computation, we have $u_{0}(p)<0$. We consider three cases.
@@
    $p>\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=0$. By Lemma 2.iii, $u_{N}(p)<0$. Thus, the unique vector $q\in Q$ satisfying $u(q,p)\geq 0$ is $q=0$. Therefore, every equilibrium induces $q=0$.
@@
    $p=\bar{p}_{N}=\bar{\nu}$. The sender-preferred no-communication outcome is $q=q^{N}$. By Lemma 2.iii, $u_{N}(p)=0$. For each $q\in Q$, we have: $u(q,p)\geq 0$ if and only if $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$, or equivalently, $q=\lambda q^{N}$ for some $\lambda\in[0,1]$. Thus, in every equilibrium, the induced buying probability is weakly smaller than in the sender-preferred no-communication outcome.
@@
    $p<\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=q^{N}$. By Lemma 2.iii, $u_{N}(p)>0$. For each $q\in Q$, we have: $u(q,p)\geq u_{N}(p)$ if and only if $q_{N}=1$ and $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$. Thus, every equilibrium induces $q=q^{N}$.

### v3

@@ 4.2 Benefiting from communication
Thus, the sender benefits from communication if and only if there exists a top equilibrium that the sender strictly prefers to the (sender-preferred) no-communication outcome. The price range in Theorem˜2 can be visualized in Figure˜1, where $N-1=3$.
@@
Next, suppose $\underaccent{\bar}{\nu}<p\leq\bar{\nu}$. Over this range, without communication, the receiver buys if and only if the checked attribute is good. Recall from Theorem˜1 that $\underaccent{\bar}{\nu}<\bar{p}_{N-1}<\bar{\nu}$. If $p\leq\bar{p}_{N-1}$, then the top-$(N-1)$ equilibrium exists. Thus, the sender strictly benefits from communication by guiding the receiver to check attributes that are better than average. If $p>\bar{p}_{N-1}$, then no top equilibrium exists. In principle, there could be other equilibria that the sender prefers to the no-communication outcome. Our proof rules this out, using the bounds in Lemma˜1.
Finally, if $p>\bar{\nu}$, then without communication, the receiver never buys. In this case, we use Lemma˜1 to show that there is no equilibrium in which the receiver buys with positive probability. The sender’s incentive constraints are critical to this conclusion. For prices $p$ slightly above $\bar{\nu}$, the receiver would be willing to buy upon seeing that one of the worst attributes was good. But in equilibrium, the sender cannot induce the receiver to check a below-average attribute. The sender would strictly prefer the receiver, while thinking he was checking one of the worst attributes, to actually check one of the best.
@@ 4.3 Optimality of top equilibria
By Theorem˜1, the top-$k$ equilibrium is the sender-preferred top equilibrium if and only if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$. Therefore, Theorem˜3 implies that the sender-preferred top equilibrium is Pareto efficient. To prove Theorem˜3, we show that if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$, then no vector of buying probabilities that satisfies the bounds in Lemma˜1 can Pareto dominate the top-$k$ equilibrium. If $p=\bar{p}_{k}$, then sender optimality follows from Pareto efficiency because the top-$k$ equilibrium gives the receiver the lowest possible payoff among all equilibria—the receiver’s payoff from the top-$k$ equilibrium is the same as his payoff from checking a random attribute and buying if and only if it is good. The prices $\bar{p}_{1},\ldots,\bar{p}_{N-1}$ can be visualized in Figure˜1 as the right endpoints of the equilibrium price intervals.
Up to this point, we have assumed that the price $p$ is exogenous. Next, we assume that the sender can choose the price (before observing the state). Thus, the sender can choose the price to maximize her revenue in the resulting equilibrium. We characterize which prices can be optimal. To state this result, it is convenient to extend our definition of the top-$k$ equilibrium to the edge cases $k=0$ and $k=N$. Under the top-$0$ strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and then always buys. Under the top-$N$ strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and buys if and only if the checked attribute is good. [^11] In each case, there are multiple strategy profiles with this property, but all such strategy profiles are payoff equivalent. Let $\bar{p}_{0}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N}=\bar{\nu}$. For $k=0$ and $k=N$, it is clear that $\bar{p}_{k}$ is the highest price at which the top-$k$ strategy profile is an equilibrium.
@@
We illustrate the intuition for Theorem˜4 graphically. Recall that $Q$ denotes the set of buying probability vectors $q=(q_{0},\ldots,q_{N})\in[0,1]^{N+1}$ satisfying the bounds in Lemma˜1. The sender’s and receiver’s payoffs from a vector $q\in Q$ are determined by two statistics: the ex-ante buying probability $x(q)$ and the expected consumption utility $y(q)$. In particular, the sender’s payoff is $px(q)$ and the receiver’s payoff is $y(q)-px(q)$. To compute these statistics, we introduce additional notation. For $j=0,\ldots,N$, let $\pi_{j}$ denote the probability $\mathbf{P}(\lvert\theta\rvert=j)$, and let $v_{j}$ denote the consumption utility $v(\theta)$ in any state $\theta$ with $\lvert\theta\rvert=j$. [^12] This is well-defined because $v$ is symmetric across the components. We have
@@
The set $\mathcal{X}$ is more convenient to analyze than the set $Q$ because it lies in $\mathbf{R}^{2}$ rather than $\mathbf{R}^{N+1}$. Figure˜2 plots the set $\mathcal{X}$ in the numerical example with $N=4$ used in Figure˜1. The upper envelope of $\mathcal{X}$ is highlighted. The extreme points of this upper envelope correspond to the top-$k$ equilibria for $k=0,\ldots,N$ and also the origin, as we show in Lemma˜3 (Section˜A.4). At price $p$, the receiver’s indifference curves in $\mathcal{X}$ are parallel lines with slope $p$. For $k=0,\ldots,N-1$, at price $\bar{p}_{k}$ the receiver is indifferent between the top-$k$ equilibrium and deviating by checking a random attribute and buying if and only if that checked attribute is good. Therefore, $\bar{p}_{k}$ is the slope of the line segment between top-$k$ equilibrium vector and the top-$N$ equilibrium vector. Using the characterization of the upper envelope of $\mathcal{X}$, we compute for each price $p$ an upper bound $V(p)$ on the sender’s revenue from all equilibria at price $p$. For $k=0,\ldots,N$, the bound $V(\bar{p}_{k})$ equals the sender’s revenue from the top-$k$ equilibrium at price $\bar{p}_{k}$. To prove Theorem˜4, it suffices to show that the function $V$ is globally maximized at some price in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. The function $V$ has kinks at these prices. We show that $V$ is convex between consecutive kinks, and hence the maximum must be achieved at one of the kinks.
The optimal price depends on the state distribution and the utility function $v$. Consider the simple example with $N=2$ and $v(\theta)=\lvert\theta\rvert$. It can be verified that the top-$1$ equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-$2$ equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. For the calculations, see Section˜A.8. When $\lvert\theta\rvert=1$, the receiver buys with certainty in the top-$1$ equilibrium but only with probability $1/2$ in the top-$2$ equilibrium. Thus, as $\pi_{1}/\pi_{2}$ increases, the buying probability ratio between the top-$1$ equilibrium and the top-$2$ equilibrium increases. The price ratio $\bar{p}_{1}/\bar{p}_{2}$ also increases because the receiver, upon seeing that a randomly chosen attribute is $1$, assigns lower probability to the event $\lvert\theta\rvert=2$.
@@ 5 Checking multiple attributes
We now extend the notion of a top-$k$ equilibrium to this setting. The top-$k$ messaging strategy for the sender is defined as in the baseline model: the sender’s message indicates the top $k$ attributes, with ties broken uniformly. The receiver’s top-$k$ strategy is as follows. The receiver checks $n\wedge k$ of the recommended attributes. Then the receiver buys if and only if the number of good attributes among those checked is at least some threshold, which is between $1$ and $n\wedge k$. [^13] Technically, we have assumed that the receiver must check $n$ attributes. Thus, the receiver can check $n-(n\wedge k)$ of the unrecommended attributes, but their values do not affect his buying probability. We say that the top-$k$ equilibrium exists if such a strategy profile is an equilibrium. [^14] In order to speak of the top-$k$ equilibrium, we assume that if there are multiple such thresholds that make the receiver’s strategy a best response, the smallest threshold is chosen.
@@
When the receiver can check multiple attributes, the set of deviations is much richer. To prove Theorem˜5, the trick is to compute the receiver’s best response in a hypothetical specification in which the receiver checks $n\wedge k$ of the $k$ recommended attributes and $n\wedge(N-k)$ of the $N-k$ unrecommended attributes. For simplicity, suppose here that $n\leq k$ and $n\leq N-k$. The random vector $(T_{n}^{k},B_{n}^{N-k})$ can only take the values
@@
We have assumed that the receiver selects all $n$ attributes before seeing any of their values. If the receiver could instead check the attributes adaptively, based on the realizations of previously checked attributes, then our proof of Theorem˜5 would still go through. Checking $n$ attributes adaptively is still less informative than checking $n\wedge k$ of the recommended attributes and $n\wedge(N-k)$ of the unrecommended attributes.
@@ A.1 Favorability
is weakly increasing in $j$ over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$; we use the convention that $a/0=\infty$ for $a>0$. If, further, the ratio in (3) is not constant over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$, then we say that $E$ is strictly more favorable than $E^{\prime}$. [^15] Our terminology is inspired by Milgrom (1981). Unlike in his definition, we do not assume that the events have full support. Our formal definition mirrors the definition of the likelihood ratio order in Shaked and Shanthikumar (2007, 1.C.1, p. 42). If $E$ is strictly more favorable than $E^{\prime}$, then $\operatorname{\mathbf{E}}[v(\theta)\mid E]>\operatorname{\mathbf{E}}[v(\theta)\mid E^{\prime}]$, by standard results about the likelihood ratio order. [^16] If $E$ is weakly more favorable than $E^{\prime}$, then the distribution of $\lvert\theta\rvert$, conditional on $E$, likelihood ratio dominates the distribution of $\lvert\theta\rvert$, conditional on $E^{\prime}$, in the sense of Shaked and Shanthikumar (2007, 1.C.1, p. 42). Moreover, likelihood ratio dominance implies first-order stochastic dominance (Shaked and Shanthikumar, 2007, Theorem 1.C.1, p. 43).
@@ A.2 Proof of Theorem 1
In Section˜A.9, we prove a more general version of Theorem˜5 that also covers the case $n=1$. Moreover, we can apply Lemma˜5 (Section˜A.9) with $k=N$ and $n=1$ to conclude that $\underaccent{\bar}{p}_{N-1,1}<\underaccent{\bar}{p}_{N,1}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N-1,1}<\bar{p}_{N,1}=\bar{\nu}$. Thus, it remains to prove that $\underaccent{\bar}{\nu}<\bar{p}_{1}$. We prove that the event $(T^{1},B^{N-1})=(1,0)$ is strictly more favorable than the event $\theta_{1}=0$; see Section˜A.1 for the definition. The event $(T^{1},B^{N-1})=(1,0)$ has support $\{1,\ldots,N-1\}$. The event $\theta_{1}=0$ has support $\{0,\ldots,N-1\}$. We have
@@ A.4 Preliminaries for Theorems 24
For $k=0,\ldots,N+1$ let $(x_{k},y_{k})=(x(q^{k}),y(q^{k}))$. Since the map $q\mapsto(x(q),y(q))$ is linear, the upper envelope of $\mathcal{X}$ can alternatively be expressed as $\cup_{k=0}^{N}[(x_{k},y_{k}),(x_{k+1},y_{k+1})]$. Figure˜2 illustrates these $N+1$ line segments in an example with $N=4$.
@@ A.5 Proof of Theorem 2
Recall from Section˜A.4 the definitions of the vectors $\Delta^{0},\ldots,\Delta^{N}$ and the functions $u_{0},\ldots,u_{N}$. We have
@@
For each $q\in Q$, each coefficient on $u_{k}(p)$ in the summation is nonnnegative. Fix a price $p$ satisfying $p>\bar{p}_{N-1}$. By Lemma˜2.iiiii, we have $u_{k}(p)<0$ for $k=1,\ldots,N-1$. By direct computation, we have $u_{0}(p)<0$. We consider three cases.
@@
    $p>\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=0$. By Lemma˜2.iii, $u_{N}(p)<0$. Thus, the unique vector $q\in Q$ satisfying $u(q,p)\geq 0$ is $q=0$. Therefore, every equilibrium induces $q=0$.
@@
    $p=\bar{p}_{N}=\bar{\nu}$. The sender-preferred no-communication outcome is $q=q^{N}$. By Lemma˜2.iii, $u_{N}(p)=0$. For each $q\in Q$, we have: $u(q,p)\geq 0$ if and only if $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$, or equivalently, $q=\lambda q^{N}$ for some $\lambda\in[0,1]$. Thus, in every equilibrium, the induced buying probability is weakly smaller than in the sender-preferred no-communication outcome.
@@
    $p<\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=q^{N}$. By Lemma˜2.iii, $u_{N}(p)>0$. For each $q\in Q$, we have: $u(q,p)\geq u_{N}(p)$ if and only if $q_{N}=1$ and $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$. Thus, every equilibrium induces $q=q^{N}$.